OpenStudy (irelandrobledo):

Which values are solutions to k-3 > –2? Check all that apply. 4

OpenStudy (danieldjpon3):

What?

OpenStudy (danieldjpon3):

k-3 > -2?

OpenStudy (danieldjpon3):

Is that the equation?

OpenStudy (irelandrobledo):

i tried doing it but it wont let me do the line but its; k-3 over 4 then >-2

alones (alones):

Waitt what about 4? is it a fraction?

OpenStudy (irelandrobledo):

yes

alones (alones):

Oh so you mean this? \[k-\frac{ 3 }{ 4 } > –2\]

OpenStudy (danieldjpon3):

Oh so 3/4

OpenStudy (danieldjpon3):

Ye that makes WAY more sense

alones (alones):

Lol.

OpenStudy (irelandrobledo):

yea sorry lol

OpenStudy (danieldjpon3):

lol

alones (alones):

So as you know you suppose to simplify those sides \[k+−\frac{ 3 }{ 4 }>−2\] and add \(\frac{ 3 }{ 4 }\) to both sides \[k+\frac{ -3 }{ 4 }+\frac{ 3 }{ 4 }>−2+\frac{ 3 }{ 4 }\]

alones (alones):

@irelandrobledo tell me what you get, and I'll check it (:<

OpenStudy (irelandrobledo):

isnt it k-3 ?

OpenStudy (irelandrobledo):

and okay thankyou ill let you know

alones (alones):

Where? x.x As your answer?

OpenStudy (irelandrobledo):

not my answer the way the problem is set up

OpenStudy (irelandrobledo):

nevermind ....but do you add 2 and 3/4

alones (alones):

Bruhh..

OpenStudy (irelandrobledo):

im sorry i really suck at math

OpenStudy (irelandrobledo):

lmfao

alones (alones):

So your answer would be \[k>\frac{ -5 }{ 4 }\]

OpenStudy (irelandrobledo):

how can you make itt into a whole number

alones (alones):

Boi, srsly x'D